OpenStudy (anonymous):

For the given statement Pn, write the statements P1, Pk, and Pk+1. 2 + 4 + 6 + . . . + 2n = n(n+1)

OpenStudy (anonymous):

i already have P_n, and P_k, i need help with P_k+1

Nnesha (nnesha):

substitute n for k+1

OpenStudy (anonymous):

okay so would it be 2k + 2(k+1) = k+1[(k+1) + 1] ?

Nnesha (nnesha):

hmm no remember on the last post \[\rm 2+4+6+......+2k=\color{ReD}{k(k+1)}\] we assumed n=k you should use this assumption for the 3rd step

Nnesha (nnesha):

http://prntscr.com/91xfp6

Nnesha (nnesha):

\[\rm \color{orange}{2+4+6+......+2k}=\color{ReD}{k(k+1)}\] n=k+1 \[\rm 2+4+6+......+2k +2(k+1)=\color{ReD}{(k+1)[(k+1)+1]}\] as you can see `2+4+6....2k ` is equal to k(k+1)

Nnesha (nnesha):

we should keep this 2+4+6....2k at left side

OpenStudy (dayakar):

p(k):2+4+ +2k = k(k+1) p(k+1): 2+4+6-----+2k+ (2k+2)= k(k+1)+(2k+2)

OpenStudy (dayakar):

is it clear

OpenStudy (anonymous):

is it k(k+1) + 2k + 2(k+1) = k+1[(k+1)+1] ?

Nnesha (nnesha):

\[\rm \color{orange}{2+4+6+......+2k}=\color{ReD}{k(k+1)}\] left side 2+4+....+2k(including) equal to k(k+1) \[\rm \color{orange}{2+4+6+......+2k} +2(k+1)=\color{ReD}{(k+1)[(k+1)+1]}\] so replace all the orange by k(k+1)

Nnesha (nnesha):